What is the Lewis Structure for CH2O ?
Answer #1

image from custom entry tool

Its two lone pairs on the oxygen (:O:) which is double bonded. Then the C is connected to the H's
Answer #2

420 H:C:H Pormal charge = - Number of bonds Number of Valence e- Number of unshaired e-s FCO) = 6-4-2=0 F.C (c) = 4 0 -4=0 FC